OpenStudy (waheguru):

plz help me in math in this question 5a+3x 3b+2x 6a-b ------ + -------- - ------- 4 3 2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{3(5a+3x)+4(3b+2x)}{12}-\frac{6a-b}{2}\] \[\frac{15a+9x+12b+8x}{12}-\frac{6a-b}{2}\] \[\frac{15a+12b+17x}{12}-\frac{6a-b}{2}\] \[\frac{15a+12b+17x-6(6a-b)}{12}\] \[\frac{15a+12b+17x-36a+6b}{12}\] \[\frac{-21a+18b+17x}{12}\]